# Dynamic Pyramid Network for Efficient Multimodal Large Language Model

### v2

@@ 1 Introduction
Despite the efficiency, we identify that these methods are sub-optimal and often sacrifice the visual representation. Specifically, most existing visual projectors are typically placed outside the LLM and downsample visual features via their saliency [10, 22]. Such straightforward practice ignores the visual semantics and the textual prompt, thus easily destroying the fine-grained visual content, e.g., OCR words and small objects. More importantly, all samples are aligned with the same visual compression rate, e.g., 75% for 2$\times$2 pooling, which inevitably hurts the understanding of hard samples and leads to the redundant computation on easy samples.
@@
To validate our approach, we adopt DPN to two popular MLLMs, i.e., LLaVA [29] and LLaVA-HR [33], and conduct extensive experiments on 10 MLLM benchmarks, e.g., MM-Vet [50]. Experiments not only confirm the great efficiency of DPN against existing methods, but also validate its competitive performance compared to the baseline. For example, compared to LLaVA-HR-X [33], our DPN-LLaVA-HR-X achieves 1.4$\times$ speedup and +0.62% performance gains. In conclusion, our contributions can be summarized in three folds:
@@ 3 Preliminaries
Plain MLLMs. As shown in Fig 1, existing MLLMs typically adopt the ViT-MLP-LLM structure, e.g., LLaVA [29]. In particular, given the input image $I\in\mathbb{R}^{H\times W\times 3}$, a ViT as visual encoder is used to extract the visual features, denoted by $X\in\mathbb{R}^{(h\times w)\times d}$. After that, the visual features are projected into the embedding space of the LLM via an MLP layer. Finally, the visual features $X\in\mathbb{R}^{(h\times w)\times d}$ are combined with the text features $Y\in\mathbb{R}^{(t-1)\times d}$ and fed into the LLM to decode the words. Therefore, a typical MLLM can be formulated by:
@@
Where $y\in Y$ and $p_{t}\in\mathbb{R}^{m}$ denote the text features and the next-token probability, respectively. $\mathcal{F}_{\text{llm}}$ and $\mathcal{F}_{\text{proj}}$ denote the LLM and the projector, respectively. $\theta$ and $\theta_{p}$ are their corresponding parameters. In Eq. 1, we can see that the computation cost of the MLLM is highly dependent on the number of the input features, i.e., $[F_{v},F_{t}]$. In practice, the number of visual features is much longer than that of textual features, e.g., 576 vs. $\sim$30 in LLaVA [29].
@@ 4.1 Overview
Here, MHA, FFN, and Norm denote the multi-head attention [42], the feed-forward network [42] and the normalization [41] , respectively. []delimited-[][$\cdot$][ ⋅ ] is the concatenation operation. $\mathcal{F}_{\text{dpe}}$ and $\theta_{d}$ denote the dynamic pooling expert layer, which will dynamically select a pooling kernel from $\{k_{0},k_{1},k_{2}\}$ to down-sample the visual features. As shown in Fig. 2, we will replace the 8-th, 16-th and 24-th layers of LLM with our DPN layer, yielding a novel dynamic pyramid structure. Compared with previous studies, such a structure not only prevents fine-grained visual contents in shallow layers, but also maximizes efficiency according to the input samples.
@@ 4.3 Routing Loss
where $C_{i}$ denotes the compression rate of the i-th pooling expert, e.g., 0.75 for 2$\times$ 2 pooling. $t$ is the target compression rate, which is set to 0.84 in our experiments.
@@ 4.5 Efficiency Analysis
Methods Params GQA SQAI VQAT VQAv2 MME MM-Vet MMB MMMU SEEDI POPE AVG FLOPs Speed
I-80B [18] 65B 45.2 - - 60.0 - - 54.5 - - - - - -
@@
Table 1: Comparison with existing efficient MLLMs. “AVG” denotes the average performance of all benchmark datasets. The best performance valuale is highlighted in bold. The speed is the average number of test samples per second and FLOPs is measured in billions.
@@ 5.2 Implementation Details
dynamic $2\times 2$,4×$\times 4$ 4 pooling 8,16,24 60.72 54.94 62.33 86.4 34.6 59.3 59.72
dynamic $1\times 2$,2×$\times 2$ 2 pooling 4,8,12 59.86 53.92 63.02 85.7 34.4 60.2 59.52
dynamic $1\times 2$,2×$\times 2$ 2 pooling 8,16,24 63.17 57.93 68.36 87.7 36.8 65.8 63.29
@@
Methods GQA VQAT POPE MMMU
Metrics FLOPs Ratio Metrics FLOPs Ratio Metrics FLOPs Ratio Metrics FLOPs Ratio
Baseline 62 100% 58.2 100% 85.9 100% 35.2 100%
+ DPE 63.17 50% 58.4 47% 87.9 48.4% 35.1 40%
+ Routing Loss 63.17 39% 57.93 36% 87.7 35% 36.8 44%
Table 4: Ablation Study of DPN on LLaVA. Our baseline is the original LLaVA without any modification.
@@ 5.4 Comparison with Existing Methods
In Figure  4, we further perform a visual analysis comparison with relevant methods to demonstrate that our approach can adjust visual token allocation based on image complexity and task specificity, highlighting its superiority in fine-grained understanding tasks. For instance, in the chart comprehension task shown in the first row of Figure b, existing methods exhibit noticeable errors. When asked, ”What is the second step suggested?”, both Fast V and TokenPacker provided incorrect answers, stating ”create list of the processes”, whereas our model correctly understood and captured the right response: ”identify your audience”. Similarly, in the text recognition task depicted in the second row, other methods also performed poorly, either missing or misidentifying certain characters, while our approach produced accurate results. Additionally, existing methods made errors in the counting task, as seen in the second image of the third row, where six birds were mistakenly identified as five. These errors highlight that direct visual compression or indiscriminate token removal can lead to significant inaccuracies. In contrast, the DPN we propose applies progressive dynamic compression in MLLMs, better preserving their ability to perceive image details.
@@ 5.4.1 Quantitative Analysis
The Diversity of Routing. We conducted a statistical analysis of the expert selection process for DPN-LLaVA 1.5 7B across different datasets, as presented in Fig 3. It is important to note that each sample undergoes expert selection three times. For simplicity, we aggregate these selections without distinguishing between layers. Generally, each dataset results in 4 to 7 distinct routing patterns.
The Impact of Image Complexity on Routing. Next, we analyzed the routing tendencies of DPE across various VQA samples, sampling from different routing patterns and visualizing the results in Figure 4 . Our observations reveal that for complex images, DPE tends to avoid pooling. For instance, in the first row, the first image features a variety of fruits and text, prompting minimal pooling. In contrast, for simpler images containing only one or a few primary objects, DPE typically applies a $1\times 2$ pooling, reducing the visual tokens by $50\%$, as shown in the second picture and third picture in the first row. For even simpler images, such as charts, DPE often utilizes $2\times 2$ pooling, reducing $75\%$ of the visual tokens, as illustrated in the fourth and fifth rows. This behavior is intuitive, as charts contain more structured and densely packed information, making them easier to represent with fewer tokens than photographs.
The Impact of Question Complexity on Routing. Finally, we perform simpler visual tasks on the first and second images of Figure 4 to explore the impact of visual tasks on routing. The results indicate that DPN can recognize simpler visual tasks and perform inference more efficiently.
For the complex fruit image in the first column, the original question likely prompted the model to focus on intricate details, such as identifying specific types of fruit or distinguishing between similarly colored objects. However, when we simplified the question to ”Is there any fruit?” the model no longer needed to process such fine-grained information. Instead, it relied on broader, more generalized visual features to determine the presence of fruit. In this case, the DPE model selected a larger kernel size, which allowed it to process larger regions of the image at once. This decision not only reduced the computational load but also improved processing speed, as larger kernels aggregate information more efficiently over wider areas.
In the case of the second image, the original reasoning-based question required extensive contextual information—such as the mood conveyed by the photo, the environment, and the clothing and expressions of the individuals. All of these factors influenced the model’s reasoning process. We switched to a simpler counting task, ”How many main characters are there in the picture?”, for testing. In this setup, DPN applies a 50% compression in the first and second DPELayer, achieving greater efficiency.

### v3

@@ 1 Introduction
Despite the efficiency, we identify that these methods are sub-optimal and often sacrifice the visual representation. Specifically, most existing visual projectors are typically placed outside the LLM and downsample visual features via their saliency [10, 21]. Such straightforward practice ignores the visual semantics and the textual prompt, thus easily destroying the fine-grained visual content, e.g., OCR words and small objects. More importantly, all samples are aligned with the same visual compression rate, e.g., 75% for $\times2$ pooling, which inevitably hurts the understanding of hard samples and leads to the redundant computation on easy samples.
@@
[…]
@@ 3 Preliminaries
Plain MLLMs. As shown in Fig. 1, existing MLLMs typically adopt the ViT-MLP-LLM structure, e.g., LLaVA [28]. In particular, given the input image $I\in\mathbb{R}^{H\times W\times 3}$, a ViT as visual encoder is used to extract the visual features, denoted by $X\in\mathbb{R}^{(h\times w)\times d}$. After that, the visual features are projected into the embedding space of the LLM via an MLP layer. Finally, the visual features $X\in\mathbb{R}^{(h\times w)\times d}$ are combined with the text features $Y\in\mathbb{R}^{(t-1)\times d}$ and fed into the LLM to decode the words. Therefore, a typical MLLM can be formulated by:
@@
Where $y\in Y$ and $p_{t}\in\mathbb{R}^{m}$ denote the text features and the next-token probability, respectively. $\mathcal{F}_{\text{llm}}$ and $\mathcal{F}_{\text{proj}}$ denote the LLM and the projector, respectively. $\theta$ and $\theta_{p}$ are their corresponding parameters. In Eq. 1, we can see that the computation cost of the MLLM is highly dependent on the number of the input features, i.e., $F_{v},F_{t}]$. In practice, the number of visual features is much longer than that of textual features, e.g., 576 vs. $\sim$30 in LLaVA [28].
@@ 4.1 Overview
Here, MHA, FFN, and Norm denote the multi-head attention [41], the feed-forward network [41] and the normalization [40] , respectively. $\cdot$] is the concatenation operation. $\mathcal{F}_{\text{dpe}}$ and $\theta_{d}$ denote the dynamic pooling expert layer, which will dynamically select a pooling kernel from $\{k_{0},k_{1},k_{2}\}$ to down-sample the visual features. As shown in Fig. 2, we will replace the 8-th, 16-th and 24-th layers of LLM with our DPN layer, yielding a novel dynamic pyramid structure. Compared with previous studies, such a structure not only prevents fine-grained visual contents in shallow layers, but also maximizes efficiency according to the input samples.
@@ 4.3 Routing Loss
where $C_{i}$ denotes the compression rate of the i-th pooling expert, e.g., 2 for $\times 2$ pooling. $t$ is the target compression rate, which is set to 1.5 in our experiments.
@@ 4.5 Efficiency Analysis
Methods Params GQA SQAI VQAT VQAv2 MME $M$-Vet MMB MMMU SEEDI POPE AVG FLOPs Speed
I-80B [laurençon2023obelicsopenwebscalefiltered] 65B 45.2 - - 60.0 - - 54.5 - - - - - -
@@
Table 1: Comparison with existing efficient MLLMs. “AVG” denotes the average performance of eight public datasets. The best performance is highlighted in bold. The speed is the average number of test samples per second and FLOPs is measured in billions.
@@ 5.2 Implementation Details
dynamic $2\times 2$$4\times 4$ pooling 8,16,24 60.72 54.94 62.33 86.4 34.6 59.3 59.72
dynamic $1\times 2$$2\times 2$ pooling 4,8,12 59.86 53.92 63.02 85.7 34.4 60.2 59.52
dynamic $1\times 2$$2\times 2$ pooling 8,16,24 63.17 57.93 68.36 87.7 36.8 65.8 63.29
@@
Methods GQA VQAT POPE MMMU Average
Metrics FR Metrics FR Metrics FR Metrics FR Metrics FR
Baseline 62 100% 58.2 100% 85.9 100% 35.2 100% 60.3 100%
+ DPE 63.17 50% 58.4 47% 87.9 48.4% 35.1 40% 61.1 46%
+ Routing Loss 63.17 39% 57.93 36% 87.7 35% 36.8 44% 61.4 38.5%
Table 4: Ablation Study of DPN on LLaVA. Our baseline is the original LLaVA without any modification. “FR” denotes the FLOPs ratio.
@@ 5.4 Comparison with Existing Methods
In Fig. 4, we further perform a visual analysis comparison with relevant methods to demonstrate that our approach can adjust visual token allocation based on image complexity and task specificity, highlighting its superiority in fine-grained understanding tasks. For instance, in the chart comprehension task shown in the first row of Fig. b, existing methods exhibit noticeable errors. When asked, “What is the second step suggested?”, both Fast V and TokenPacker provided incorrect answers, stating “create list of the processes”, whereas our model correctly understood and captured the right response: “identify your audience”. Similarly, in the text recognition task depicted in the second row, other methods also performed poorly, either missing or misidentifying certain characters, while our approach produced accurate results. Additionally, existing methods made errors in the counting task, as seen in the second image of the third row, where six birds were mistakenly identified as five. These errors highlight that direct visual compression or indiscriminate token removal can lead to significant inaccuracies. In contrast, the DPN we propose applies progressive dynamic compression in MLLMs, better preserving their ability to perceive image details.
@@ 5.4.1 Quantitative Analysis
The Diversity of Routing. We conducted a statistical analysis of the expert selection process for DPN-LLaVA 1.5 7B across different datasets, as presented in Fig. 3. It is important to note that each sample undergoes expert selection three times. For simplicity, we aggregate these selections without distinguishing between layers. Generally, each dataset results in 4 to 7 distinct routing patterns.
The Impact of Image Complexity on Routing. Next, we analyzed the routing tendencies of DPE across various VQA samples, sampling from different routing patterns and visualizing the results in Fig. 4 . Our observations reveal that for complex images, DPE tends to avoid pooling. For instance, in the first row, the first image features a variety of fruits and text, prompting minimal pooling. In contrast, for simpler images containing only one or a few primary objects, DPE typically applies a $1\times 2$ pooling, reducing the visual tokens by $50\%$, as shown in the second picture and third picture in the first row. For even simpler images, such as charts, DPE often utilizes $2\times 2$ pooling, reducing $75\%$ of the visual tokens, as illustrated in the fourth and fifth rows. This behavior is intuitive, as charts contain more structured and densely packed information, making them easier to represent with fewer tokens than photographs.
The Impact of Question Complexity on Routing. Finally, we perform simpler visual tasks on the first and second images of Fig. 4 to explore the impact of visual tasks on routing. The results indicate that DPN can recognize simpler visual tasks and perform inference more efficiently.
For the complex fruit image in the first column, the original question likely prompted the model to focus on intricate details, such as identifying specific types of fruit or distinguishing between similarly colored objects. However, when we simplified the question to “Is there any fruit?” the model no longer needed to process such fine-grained information. Instead, it relied on broader, more generalized visual features to determine the presence of fruit. In this case, the DPE model selected a larger kernel size, which allowed it to process larger regions of the image at once. This decision not only reduced the computational load but also improved processing speed, as larger kernels aggregate information more efficiently over wider areas. In the case of the second image, the original reasoning-based question required extensive contextual information—such as the mood conveyed by the photo, the environment, and the clothing and expressions of the individuals. All of these factors influenced the model’s reasoning process. We switched to a simpler counting task, “How many main characters are there in the picture?”, for testing. In this setup, DPN applies a 50% compression in the first and second DPE layer, achieving greater efficiency.